Respected sir,
[1] Rs.1980 are divided among A,B,C so that half of A's part ,one-third of B's part and one-sixth of C's part are equal.Then , B's part is ?

plz help me to solve ratio ,proportions maths problem.waiitng for yur reply.
thank u sir.
regards.
Binita

Answer by RAY100(1637)

1980=A+B+C
and A/2=B/3=C/6
mult by 6
3A=2B=C
or A=C/3,,,,and B=C/2
subst
1980=C/3+C/2+C
1980=(2C+3C+6C)/6
1980=(11/6)C
1080=C
and therefore
A=C/3=360
B=C/2=540
check
1980=A+B+C=360+540+1080=1980,,,ok